Corporate Finance
(15 points) Suppose the risk-free rate is 6.3% and the market portfolio has an expected rate of return of 14.8%. The market portfolio has a variance of 0.0121. Portfolio Z has a correlation coefficient with the market of 0.45 and a variance of 0.0169. According to CAPM, what is the expected rate of return on portfolio Z? 4.
Answer #1

For the above question,

  • Risk free rate ( r​​​​​​ f ) = 6.3%
  • Expected rate of return E(rm) = 14.8%
  • Variance ( \sigma2m )= 0.0121
  • Portfolio Z correlation coefficient ( \rhoz,m)= 0.45
  • Portfolio Z variance = 0.0169

We know, \rho​​​ ​​​​​​​​z,m = Cov (Rz, Rm) / (0.02)(0.57)   \therefore Cov (Rz, Rm) = 0.03192

Also, \betaz = Cov (R​​​​​​ z , R m ) / \sigma2m = 0.03192 / 0.0121 (Had Put the value) = 2.6380

Thus, Expected rate of return on portfolio Z :-

= 6.3% + 2.6380*( 14.8% - 6.3% ) = 2.87%  

[ Note : Expected rate of return = Risk free rate + beta*(market return rate - risk free rate) ]

  

​​​